OpenStudy (anonymous):

in triangle XYZ, x=20.4, y=16.4, and z=23.1, find the m

OpenStudy (anonymous):



OpenStudy (adi3):

triangles

OpenStudy (adi3):

@lyssaaaa

OpenStudy (adi3):

First we need to remember that all angles of triangles = 180

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thats it? @Adi3

OpenStudy (adi3):

no

OpenStudy (adi3):

if i do it in my way i would cross multiply

OpenStudy (adi3):

look at this. I lost connection @lyssaaaa

OpenStudy (adi3):

do u get it. Quick

OpenStudy (anonymous):

then is it transfer to the other side or divide

OpenStudy (anonymous):

**after the multiplying

OpenStudy (adi3):

20.4 * 16.4 = 334.56 23.1z z = 334.56/23.1 =

OpenStudy (adi3):

do u get it quick

OpenStudy (anonymous):

yes thank you

OpenStudy (adi3):

check ur answer first maybe correct i'm not sure

OpenStudy (adi3):

because ITS FIND ANGLE Z not SIDE

OpenStudy (adi3):

@lyssaaaa

OpenStudy (anonymous):

i get it but is der lyk cos/sin/tan bec we're kinda lyk in the law of cosines idk

OpenStudy (adi3):

ohh I'm just in 10 and till yet we did not get cos, sin, tan we will get in 11

OpenStudy (anonymous):

but im only year 9 omg

OpenStudy (adi3):

actually I lied I'm also in 9th

OpenStudy (anonymous):

haha

OpenStudy (adi3):

but my school is a american school and the rule is that we will learn sin, cos, tan in 10

OpenStudy (adi3):

wht school do u go to @lyssaaaa

OpenStudy (anonymous):

ohhh im not american thats why :((

OpenStudy (anonymous):

thats lyk so unfair im already dying here omg trigonometry pls

OpenStudy (adi3):

wht r u

OpenStudy (anonymous):

i am human

OpenStudy (adi3):

no u r a alien

OpenStudy (anonymous):

yes i live in space

OpenStudy (adi3):

no in black hole

OpenStudy (anonymous):

u evil

OpenStudy (anonymous):

brb

OpenStudy (adi3):

wht u said u r an human even a 1 year baby can tell u r a human. Until u look is so bad that u literally look like a alien

OpenStudy (adi3):

do look that bad

OpenStudy (adi3):

from ur name @lyssaaaa I cna tell that u r indian

OpenStudy (adi3):

replyyy wht happened

OpenStudy (adi3):

@lyssaaaa

OpenStudy (adi3):

guys can u give me a good topic for child soldier I WILL MEDAL THAT PERSON

OpenStudy (adi3):

@lyssaaaa

OpenStudy (adi3):

@Compassionate

OpenStudy (adi3):

@munna14

OpenStudy (anonymous):

hiii lol

OpenStudy (adi3):

guys can u give me a good topic for child soldier

OpenStudy (adi3):

quick

OpenStudy (adi3):

@lyssaaaa u still din't close it